Fifteen items or less: The number of customers in line at a supermarket express checkout counter is a random variable with the following probability distribution.

x

0

1

2

3

4

5

Px

0.05

0.25

0.35

0.20

0.10

0.05

Compute the standard deviation σX. Round the answer to at least three decimal places.
Answer #1

Solutionj:

x P(x) x * P(x) x2 * P(x)
0 0.05 0 0
1 0.25 0.25 0.25
2 0.35 0.7 1.4
3 0.2 0.6 1.8
4 0.1 0.4 1.6
5 0.05 0.25 1.25
Sum 1 2.2 6.3

Mean = X * P(X)

= 2.2

Standard deviation =

=X 2 * P(X) - 2

=  6.3 - (2.2)2 = 1.46

Standard deviation = = 1.208